OpenStudy (anonymous):

Jacqui has grades of 78 and 87 on her first two algebra tests. If she wants an avergae of at least 76, what possible scores can she make on her third?

OpenStudy (anonymous):

76 >= (78 + 87 + t)/3 228 >= 78 + 87 + t 228 >= 165 + t 63 >= t SO she must get a 63 or above

OpenStudy (anonymous):

I think the > is in the wrong direction but the answer is right

OpenStudy (amistre64):

\[\frac{\sum(priors)+new}{\#priors+1}\ge76\] \[\sum(priors)+new\ge76(\#priors+1)\] \[new\ge76(\#priors+1)-\sum(priors)\] maybe

OpenStudy (amistre64):

this is greater than or equal to that ... it fits

OpenStudy (anonymous):

Great thanks a lot:)

OpenStudy (anonymous):

sorry I am a bit confused so she has to get 63 or greater or greater than 63

OpenStudy (amistre64):

yes, they have to make at least a 63 to average out to a 76. anything less than that will not do.